A triangle has vertices K(-2,2), L(1,5) and M (3,-3). Verify that the triangle has a right angle

Answer by J2R2R(94)

A triangle has vertices K(-2,2), L(1,5) and M (3,-3). Verify that the triangle has a right angle.

Gradient of KL = (5-2)/(1--2) = 3/3 = 1
Gradient of LM = (-3-5)/(3-1) = (-8)/2 = -4
Gradient of MK = (2--3)/(-2-3) = 5/(-5) = -1

A right angle means that the product of the gradients of the lines making the right angle is -1.
Here we see KL and MK are at right angles since the product of their gradients is -1. Thus K must be a right angle.